Please help me solve the following problem: A survey of 1000 NUST students find that 48% own a personal computer. Construct a 95% confidence interval on the true proportion of NUST students who own a personal computer.
Answer by rothauserc(4718)

Our sample size of 1000 is sufficiently large > 30 that we can use the normal distribution
:
sample proportion(p) = 0.48
:
standard error(SE) = square root(0.48 * (1-0.48) / 1000) = 0.0158
:
alpha(a) = 1 - (95/100) = 0.05
:
critical probability(p*) = 1 - (a/2) = 0.975
:
critical value(CV) is the z-value associated with p* which is 1.96
:
margin of error(ME) = CV * SE = 1.96 * 0.0158 = 0.031
:
**************************************************************
95% confidence interval is 0.48 + or - 0.031 or (0.449, 0.511)
**************************************************************
: